A magazine collects data each year on the price of a hamburger in a certain fast food restaurant in various countries around the world. The price of this hamburger for a sample of restaurants in Europe in January resulted in the following hamburger prices (after conversion to U.S. dollars).

5.19 4.95 4.07 4.69 5.23 4.67
4.12 4.99 5.15 5.55 5.34 4.60

The mean price of this hamburger in the U.S. in January was $4.62. For purposes of this exercise, assume it is reasonable to regard the sample as representative of these European restaurants. Does the sample provide convincing evidence that the mean January price of this hamburger in Europe is greater than the reported U.S. price? Test the relevant hypotheses using

α = 0.05.

(Use a statistical computer package to calculate the P-value. Round your test statistic to two decimal places and your P-value to three decimal places.

t =

P-value

= .351

State your conclusion.

Reject H0. We have convincing evidence that the mean price of a hamburger in a certain fast food restaurant in Europe is greater than $4.62.

I am missing T
Answer #1

Ans:

sample size,n=12

sample mean=4.879

sample standard deviation=0.4633

Test statistic:

t=(4.879-4.62)/(0.4633/SQRT(12))

t=1.94

df=12-1=11

p-value=tdist(1.9365,11,1)=0.039

Reject H0. We have convincing evidence that the mean price of a hamburger in a certain fast food restaurant in Europe is greater than $4.62.